# post-last exit time process and its application to Loss-Given-Default Distribution

### v3

@@ 3.2. Loss-given-default distribution
$\mbox{Erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}{\rm d}t\quad\text{and}\quad\mbox{Erfc}(x)=1-\mbox{Erf}(x).$
As announced above (right after the beginning of this subsection), we shall find $\alpha^{*}$ that matches (3.11) with $T=5$ with the market-quoted 5-year default probability. $\diamondsuit$
@@
[…]
@@
Define $Z^{*}_{t}:=\alpha^{*}-Z_{t}$. Then, by (3.3)
where $(W^{*}_{t})_{t\geq 0}:=(-W_{t})_{t\geq 0}$ is a standard Brownian motion. We have used the fact that $-M\coth(-M\cdot x)=M\coth(M\cdot x)$. We have $\tau=\min\left(s:\int_{0}^{s}\mbox{1}\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)=\min\left(s:\int_{0}^{s}\mbox{1}\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)$ and
$\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ $=\lim_{z^{\prime}\uparrow\alpha^{*}}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\mid Z_{0}=z^{\prime}\right]$
$=\lim_{z\downarrow 0}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right].$ (3.13)
By this translation to $Z^{*}$, (3.8) becomes $U(z)=\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right]$ and (3.9) is now
$\frac{1}{2}U^{\prime\prime}(z)+|M|\coth(|M|\cdot z)U^{\prime}(z)-\left(1+\gamma\right)U(z)$ $=-\mbox{1}\mbox{l}_{[\alpha^{*}-Q,+\infty)}(z),\quad z\in(0,\infty).$ (3.14)
(i) Consider the case when $z\geq\alpha^{*}-Q$. Then (3.14) is reduced to
$\psi_{1}(z)=\frac{\sqrt{2}\sinh\left(|M|b_{1}z\right)}{\sqrt{\pi}b_{1}\sinh(|M|\cdot z)},\quad\phi_{1}(z)=\frac{\sqrt{\pi}e^{-|M|b_{1}z}}{\sqrt{2}\sinh(|M|\cdot z)},$ (3.15)
(ii) Let $0<z<\alpha^{*}-Q$. In this case, (3.14) is reduced to
As in the case (i) above, its fundamental solutions are given by $\psi_{1}(z)$ and $\phi_{1}(z)$ in (3.15). Therefore, in this case we have $U(z)=d\psi_{1}(z)$ with some constant $d$. We have eliminated the second fundamental solution because $U(z)$ is bounded when $z\to 0$.
Finally, $\lim_{z\downarrow 0}U(z)=\lim_{z\downarrow 0}d\psi_{1}(z)=d\sqrt{\frac{2}{\pi}}$, which proves (3.12) in view of (3.2). ∎
@@
Proof.
$\mathbb{P}_{\alpha^{*}}\left(Z_{\tau}\leq Q\right)=\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\mbox{l}_{\{Z_{\tau}\leq Q\}}\right].$
Thus, we obtain the desired result by taking the limit $\gamma\downarrow 0$ in (3.12). ∎
@@
Proof.
[…]
Then, the density of the transform $1-e^{\sigma Z_{\tau}}$ is given by
Finally, $\int_{1-e^{\sigma\alpha^{*}}}^{1}x\mathbb{P}_{\alpha^{*}}(K^{B}(\xi)\in{\rm d}x)$ provides (3.17). ∎
@@ 3.3. Parameter estimation
Fix any point in time as the current time. The procedure outlined below estimates the LGD distribution implied in the 5-year CDS spread using the information available up to the current time. The parameters are estimated in a way that the default probability and default time distribution obtained from the model are consistent with the default probability provided by Eikon (consistent with the quoted CDS spread).
@@
    Estimate the parameters $(\mu,\sigma)$ of the asset process in (3.1). Since it is a well-known procedure, detailed explanation is delegated to Appendix A.2. The initial value $Y_{0}$ of the leverage ratio process is retrieved using the estimated $\sigma$. The debt process $B$ represents the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach): the reason being due to the good fit to empirical evidence.
@@
    Calibrate $\alpha$ so that the default time $\xi$ in (3.6) matches the market-quoted 5-year default probability by using the method described in Remark 3.1, especially (3.11).
@@
    We obtain the LGD distribution of $K^{B}(\xi)$ from Corollary 3.2. However, in the case of CDS, it is more appropriate to consider the LGD distribution of the total debt $D=(D_{t})_{t\geq 0}$, which is done as follows:
@@ 3.4. Model Check
We set the principal amount to \$$\$1$ and assume the spread payments are made quarterly. We find spread value for which the present value of spread payments (premium leg) and the present value of the payment in case of default (default leg) are equal. We ignore the counterparty risk and use the constant rate $r$ from Step 1 as the discount rate. The following procedure is for estimating $\rho$.
@@
    We shall obtain samples of $\xi(\omega_{i})$ by using $\tau(\omega_{i})$ in Step (i). Simulate $L_{\alpha}(\omega_{i})$ by inverting its cumulative distribution function (which is also used in (3.11)) at $U^{\prime}(\omega_{i})$ (uniformly distributed random variable on $[0,1]$, independent of $U(\omega_{i})$). We thus obtain $\xi(\omega_{i})=L_{\alpha}(\omega_{i})+\tau(\omega_{i})$.
@@
    We have obtained the pairs $(\xi(\omega_{i}),K^{D}(\xi(\omega_{i}))$ from which we can compute the present value of the recovery (on the dollar) upon default. Specifically, we calculate the value of $e^{-r(\xi(\omega_{i}))}K^{D}(\xi(\omega_{i}))\mbox{1}\mbox{l}_{\{\xi(\omega_{i})\leq 5\}}$. By taking the average over $N$ trials, we find the value of default leg.
@@
    Using the discount rate $r$, calculate the present value of spread payments until $\xi(\omega_{i})$ (including accrued premium) or the maturity, whichever is earlier. By taking average over $N$ trials, we obtain the premium leg.
@@
4. Example of Tyson Foods Inc.
In this section, we consider an example of Tyson Foods Inc., hereafter denoted shortly as Tyson Foods. The S&P Global credit rating of this company was downgraded from BBB+/Negative to BBB/Stable on November 28, 2023 (S&P Global, 2023). The main purpose of presenting this example is to illustrate how to implement our method to obtain the LGD distribution implied in the current credit market and company-specific financial conditions. Specifically, we estimate the model parameters in a way that the model-implied default probability (which is based on default time distribution) is consistent with the default probability provided by Eikon. Based on those parameters, we obtain the LGD distribution.
We choose December 29, 2023 as the current time and use the information available up to this time to obtain the LGD distribution implied on that day. We use 1-year daily data from the period 2022/12/29$\sim$ 2023/12/29 for the parameter estimation. We use 1-year Treasury bill rate (bank discount rate) $4.55\%$ (of December 29, 2023) as the constant rate $r$. The current quoted 5-year CDS spread for Tyson Foods is 69.34 bps and 5-year default probability is $5.965\%$. The values of the equity $E$ and debt $B$ are displayed in Figure 2. We interpolated daily debt values using standardized quarterly balance sheets. We defined the debt $B$ as the sum of the short-term debt (notes payables, short-term debt, current portion of long-term debt and capital leases) and one half of the long-term debt (long-term debt and capital leases).
[…]
Figure 2. Daily values of Tyson Foods data during 2022/12/29$\sim$ 2023/12/29: (a) Market capitalization $E$ in million USD and (b) Debt $B$ in million USD.
[…]
Table 3. Estimated parameters up to 4 decimal points. The standard error of $\sigma$ from the maximum likelihood estimation is given in parenthesis.
$\sigma$ 0.2499 (0.0112)
$\mu$0.0704
$M$0.5888
$Y_{0}$ 3.2693
$\alpha$ 0.9304
Figure 3. 5Y default probability (3.11) as a function of $\alpha$. Market-quoted 5Y default probability is $5.965\%$.
Finally, we check the consistency of the model-implied LGD distribution with the CDS market. We set the number of trials $N$ as 100,000. Following the procedure in Section 3.4, we calculate the model-implied 5-year CDS spread which is based on the model-implied LGD distribution $K^{D}(\xi)$. This spread is $57.8976$ bps. While $\mathbb{E}_{\alpha^{*}}[K^{D}(\xi)]=57.2669\%$ from (3.17) and (3.18), we shall go with the average LGD: $\mathbb{E}_{\alpha^{*}}\left[K^{D}(\xi)\mid\xi\leq 5\right]$ since we have samples of $K^{D}(\xi(\omega_{i}))$ and $\xi(\omega_{i})$ in Step (i) and Step (ii), respectively. The estimated average LGD is $51.6195\%$. In the model’s case, $\widehat{\rho}=57.8976/51.6195=1.1216$bps, while the counterpart of the market quote is $\rho_{\text{quote}}:=69.34/60=1.1557$bps. We see that $\widehat{\rho}$ and the market estimate $\rho_{\text{quote}}$ are very close. Therefore, our method is consistent with the market information.
Furthermore, to check the robustness of our results with respect to the ratio of long-term debt to total debt (denoted by $w$), we report the results for four additional values of $w$ together with $\widehat{w}=70.1037\%$ (obtained in Step 4 of Section 3.3) in Table 4. We see that all of our estimates of $\rho$ are within $3\%$ range from the market estimate of $\rho_{\text{quote}}=1.1557$bps. We also notice that these $\hat{\rho}$’s themselves are stable with respect to $w$. We conclude that the default time and LGD distribution in our framework are consistent with the credit market and company-specific financial conditions.
[…]

### v4

@@ 4.1. Loss-given-default distribution
$\mbox{Erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}{\rm d}t,\qquad\mbox{Erfc}(x)=1-\mbox{Erf}(x),$
and $G(\cdot,\cdot)$ is defined in (2.1).
Proof.
See Appendix A.5
As announced in the beginning of Section 2, we shall find $\alpha^{*}$ that matches (4.9) with $T=5$ with the market-quoted 5-year default probability.
Finally, we undertake the procedure to compute the distribution of the asset-to-debt ratio at time $\xi$ (i.e., $Y_{\xi}$), or equivalently that of its post-last passage time process at $\tau$ (i.e., $Z_{\tau}$) in view of (4.4).
@@
$\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]=\frac{1}{1+\gamma}\left(\cosh\left(|M|(\alpha^{*}-Q)\right)+b_{1}\sinh\left(|M|(\alpha^{*}-Q)\right)\right)e^{-b_{1}|M|(\alpha^{*}-Q)}.$ (4.10)
@@
The proofs of this proposition along with the subsequent corollaries are in Appendix A.6. ∎
@@
From this result we calculate the distribution of LGD $K^{B}(\xi)$ in (4.5):
@@ 4.2. Parameter estimation
Fix any point in time as the current time. The simple procedure outlined below estimates the LGD distribution implied in the 5-year CDS spread using the information available up to the current time. The parameters are estimated in a way that the default probability and default time distribution obtained from the model are consistent with the default probability provided by LSEG Workspace (consistent with the quoted CDS spread).
@@
    Estimate the parameters $(\mu,\sigma)$ of the asset process in (2.4). Since it is a well-known procedure, detailed explanation is delegated to Appendix A.7. The current value $Y_{0}$ of the asset-to-debt ratio process is retrieved using the estimated $\sigma$. The debt process $B$ represents the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach): the reason being due to the good fit to empirical evidence.
@@
    Calibrate $\alpha$ so that the default time $\xi$ in (4.4) matches the market-quoted 5-year default probability by using Proposition 4.1.
@@
    Obtain the LGD distribution of $K^{B}(\xi)$ from Corollary 4.2. However, in the case of CDS, it is more appropriate to consider the LGD distribution of the total debt $D=(D_{t})_{t\geq 0}$, which is done as follows:
@@ 4.3. Model Check
We set the principal amount to $\mathdollar 1$ and assume the spread payments are made quarterly. We find spread value for which the present value of spread payments (premium leg) and the present value of the payment in case of default (default leg) are equal. We ignore the counterparty risk and use the risk-free rate given by $r$ as the discount rate. The following procedure is for estimating $\rho$.
@@
    We shall obtain samples of $\xi(\omega_{i})$ by using $\tau(\omega_{i})$ in Step (i). Simulate $\lambda_{\alpha}(\omega_{i})$ by inverting its cumulative distribution function (which is also used in (4.9)) at $U^{\prime}(\omega_{i})$ (uniformly distributed random variable on $[0,1]$, independent of $U(\omega_{i})$). We thus obtain $\xi(\omega_{i})=\lambda_{\alpha}(\omega_{i})+\tau(\omega_{i})$.
@@
    We have obtained the pairs $(\xi(\omega_{i}),K^{D}(\xi(\omega_{i}))$ from which we can compute the present value of the recovery (on the dollar) upon default. Specifically, we calculate the value of $e^{-r\cdot\xi(\omega_{i})}K^{D}(\xi(\omega_{i}))\mbox{1}\hskip-2.5pt\mbox{l}_{\{\xi(\omega_{i})\leq 5\}}$. By taking the average over $N$ trials, we find the value of default leg.
@@
    Using the discount rate $r$, we calculate the present value of spread payments until $\xi(\omega_{i})$ (including accrued premium) or the maturity, whichever is earlier. By taking average over $N$ trials, we obtain the premium leg.
@@
5. Empirical Study
5.1. Data
In this section, we analyze eight companies listed on New York Stock Exchange (NYSE). These firms together with their NYSE ticker symbols are as follows: Delta Air Lines, Inc. (DAL), Ford Motor Company (F), Kohl’s Corporation (KSS), Lennar Corporation (LEN), Toll Brothers, Inc. (TOL), Tyson Foods, Inc. (TSN), Whirlpool Corporation (WHR) and The Williams Companies, Inc. (WMB). In the sequel, we use NYSE ticker symbols for brevity. According to the Repository Reporting OTC Data on the Depository Trust and Clearing Corporation (DTCC) website, these firms have average daily notional amount of more than 15 million USD in CDS contracts during March 22, 2025 to June 20, 2025 (Q2)11endnote: 1https://www.dtcc.com/repository-otc-data (Accessed on 16/09/2025). Hence we believe that these names have sufficient liquidity in the CDS market. The selected companies are presented in Table 2 together with their industry classifications 22endnote: 2Industry classifications can be checked on Nasdaq website https://www.nasdaq.com/market-activity/stocks/screener (Accessed on 18/09/2025). The main purpose of presenting these examples is to illustrate how to implement our method to obtain the LGD distribution implied in the current credit market together with company-specific financial conditions. Specifically, we estimate the model parameters in a way that the model-implied default probability (which is based on default time distribution) is consistent with the default probability provided by LSEG Workspace. Based on those parameters, we obtain the LGD distribution. See Steps 1 through 3 in Section 4.2.
For each company, we choose one time point as the current day and use the information available up to that time point to obtain the LGD distribution implied in the market on that time point. We use 6-month daily data for the parameter estimation. The 6-month periods selected for each company are listed in Table 2. Since the purpose is to estimate the LGD distribution of relatively high credit risk scenarios, we selected the latest available 6-month periods which yield estimated drift parameter $M$ and barrier $\alpha$ satisfying $M<0$ and $\alpha\leq 1$.
We use the market yield on U.S. Treasury Securities at 5-year constant maturity (quoted on an investment basis, not seasonally adjusted) as the constant rate $r$. These rates were obtained from the Federal Reserve Bank of St. Louis website. We interpolated daily debt values using standardized quarterly balance sheets. We defined the debt $B$ as the sum of the short-term debt (notes payables, short-term debt, current portion of long-term debt and capital leases) and one half of the long-term debt (long-term debt and capital lease obligations). The average value of the ratio of long-term debt to total debt is given by $\widehat{w}$.
5.2. Procedure
Table 2 reports the estimated parameters of the asset-to-debt ratio processes; $\sigma,\mu,M$, and implied current $Y_{0}$ up to the first eleven rows. This is done by Step 1 in Section 4.2. The level of $\alpha$ is calibrated to match model-implied 5-year probability of default (PD) with the market counterpart (obtained from LSEG Workspace). See Step 2 of Section 4.2. The reader can see that the numbers in the line headed by ‘Market PD (5Y)’ are exactly the same as those in the line headed by ‘Model-implied PD (5Y)’ across the board: there is no arbitrariness in choosing $\alpha$. The third step is to compute $K^{D}(\xi)$ by Corollary 4.1. Figure 1 displays the density function of $K^{D}(\xi)$ in (4.13).
As a final procedure, we check the consistency of the model-implied LGD distribution with the current CDS market. Refer to Table 2 again. Following the procedure in Section 4.3 and setting the number of trials $N$ to 10,000, we calculate the current model-implied 5-year CDS spread based on the model-implied LGD distribution of $K^{D}(\xi)$. See the line headed by ‘Estimated CDS spread (5Y)’. Note that in the calculation process we use samples of $K^{D}(\xi(\omega_{i}))$ and $\xi(\omega_{i})$ from Step (i) and Step (ii) of Section 4.3, respectively. For each company, we calculate model-implied $\rho$ (‘Model $\rho$’) and compare it to the market $\rho$ (‘Market $\rho$’) which is the ratio of quoted CDS spread over 60% loss rate.
Figure 2 shows the density function of $\lambda_{\alpha^{*}}$, where $\alpha^{*}=\frac{1}{\sigma}\ln(\alpha)$. Recall the normalized asset-to-debt ratio process $\frac{1}{\sigma}\ln(Y)$ in (2.5) with state space $(\ell,r)=(-\infty,+\infty)$ whose killing boundary, based on which the $\lambda_{\alpha^{*}}$ density function is computed, is the left natural boundary $\ell=$-$\infty$; no killing in the interior. In Figure 3, we provide the density of the occupancy time above $\alpha^{*}$ before killing, which is denoted by $\lambda_{\alpha^{*}}^{A,c}$. The process $\frac{1}{\sigma}\ln(Y)$ is forcibly killed on the first instant to level $c^{*}=\frac{1}{\sigma}\ln(c)$. The density of $\lambda_{\alpha^{*}}^{A,c}$ shows, based on the firm-value approach, how much time the asset-to-debt ratio stays above $\alpha^{*}$ before it hits $c^{*}$. From the viewpoint of risk management, we can set $c^{*}$ at a landmark point for decision making. In this analysis, we set $c=0$, $\frac{2}{3}\alpha$, and almost equal to $\alpha$. The case $c\rightarrow 0$ corresponds to $c^{*}\rightarrow$-$\infty$, which is one of the natural boundaries of $\frac{1}{\sigma}\ln(Y)$. By comparing $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c=0}$, we see how much time the process $\frac{1}{\sigma}\ln(Y)$ spends above level $\alpha^{*}=\frac{1}{\sigma}\ln(\alpha)$, relative to the last passage time $\lambda_{\alpha^{*}}$. See the end of this section (as well as A.7) for the computation of Figures 2 and 3.
Table 2. Estimated parameters up to 4 decimal points together with the standard error of $\ln(\sigma)$ produced by the maximum likelihood estimation. Company industry classifications in order: Meat/Poultry/Fish, Homebuilding, Natural Gas Distribution, Homebuilding, Air Freight/Delivery Services, Consumer Electronics/Appliances, Auto Manufacturing, Department/Specialty Retail Stores.
TSN TOL WMB LEN DAL WHR F KSS
Dataset start date 2024-12-11 2025-01-29 2022-11-29 2024-11-25 2024-12-24 2024-12-24 2024-12-04 2025-01-30
Dataset end date=Current Date 2025-06-16 2025-07-31 2023-06-01 2025-05-30 2025-06-30 2025-06-30 2025-06-09 2025-08-01
Current risk-free rate 0.0404 0.0396 0.037 0.0396 0.0379 0.0379 0.0409 0.0377
$\hat{w}$ 0.9317 1.0000 0.9374 1.0000 0.8376 0.7135 0.6496 0.8399
Estimated $\ln(\sigma)$ -1.6378 -1.0399 -1.6882 -1.1320 -0.8134 -1.3957 -2.3723 -0.9104
Standard Error 0.0630 0.0631 0.0630 0.0630 0.0631 0.0640 0.0241 0.1003
Log-Likelihood -903.7338 -885.8554 -980.2036 -1005.2365 -1072.2375 -812.9502 -1029.2806 -711.4242
$\sigma$ 0.1944 0.3535 0.1849 0.3224 0.4433 0.2477 0.0933 0.4024
$\mu$ -0.1728 -0.2310 -0.1741 -0.9825 -0.2969 -0.0741 -0.0117 -0.3108
$M$ -1.1937 -0.9423 -1.2344 -3.3316 -0.9769 -0.5761 -0.6106 -1.0672
Implied Current $Y_{0}$ 5.1854 8.7477 3.9325 18.0141 4.7479 2.1385 1.3728 1.3409
$\alpha$ 0.9967 0.8307 0.9215 0.0592 0.2689 0.8190 0.9598 0.3541
Expected value of $K^{D}(\xi)$ 0.6138 0.7503 0.6411 0.9867 0.9160 0.6261 0.4355 0.8855
Market PD (5Y) 4.04 6.824 8.177 8.445 9.062 13.95 16.208 58.142
Model-implied PD (5Y) 4.0400 6.8241 8.1769 8.4451 9.0621 13.9499 16.2077 58.1419
Market CDS spread (5Y) 47.29 83.14 96.05 99.52 136.68 171.51 201.4 925.32
Estimated CDS spread (5Y) 41.4160 88.6270 90.3355 152.7450 153.7701 159.6580 130.6061 1182.6249
5Y PD from CDS estimation (Step(v)) 4.04 6.82 8.17 8.44 9.06 13.96 16.2 58.16
Market $\rho$ 0.7882 1.3857 1.6008 1.6587 2.2780 2.8585 3.3567 15.4220
Model $\rho$ 0.7520 1.2829 1.5425 1.5617 1.7220 2.7382 3.1847 13.5751
(Model $\rho$ - Market $\rho$)/Market $\rho$ -0.0459 -0.0742 -0.0365 -0.0584 -0.2441 -0.0421 -0.0512 -0.1198
5.3. Analysis
The market data is consistent with the fact that higher default probability corresponds to higher CDS spread. Estimated CDS spread is based on the distribution of $K^{D}(\xi)$ which is different from constant 60% loss rate used as market convention; therefore, estimated and market spreads will naturally differ. However, as shown in the bottom row of Table 2, for six out of eight companies their percentage difference of Model and Market $\rho$’s is less than $7.42\%$. We use Table 2 together with Figure 1 for the comparison. Figure 1 suggests that for ‘TSN’, ‘WMB’, and ‘WHR’, $K^{D}(\xi)$ is distributed somewhat around 60%. These are three companies whose expected loss rate (‘Expected value of $K^{D}(\xi)$’) is closest to 60%. We look at the bottom row of Table 2 to find that these three companies have the smallest percentage difference (in absolute value) between ‘Market $\rho$’ and ‘Model $\rho$’, no more than 4.6%. Based on these observations, we can state that the model is consistent with the market. The exceptions are ‘DAL’ and ‘KSS’. For these two companies, the deviation of ‘Market $\rho$’ and ‘Model $\rho$’ is the largest (24% and 12% respectively). In the case of very high CDS spread levels (like ‘KSS’), prices are more susceptible to liquidity (supply-demand gap) than to fundamental credit quality; and hence prices become volatile and unpredictable. Hence the deviation of ‘KSS’ ‘Market $\rho$’ from ‘Model $\rho$’ is understandable. For the remaining case of ‘DAL’, it seems reasonable to attribute the discrepancy of the $\rho$’s to the high volatility $\sigma=0.4433$, which may contribute to the difficulty in pricing spreads. Moreover, we point out an observation common to ‘DAL’ and ‘KSS’: the distribution of their $K^{D}(\xi)$ is also concentrated at much higher levels (around 90%) than 60%. From these, we derive a conclusion that the closer the distribution of $K^{D}(\xi)$ is to 60% loss rate, the estimated spread per 1% loss rate (which is the definition of $\rho$) is sufficiently close to the market counterpart. On the other hand, this conclusion suggests that for companies whose loss rate distribution is concentrated in a region far from 60%, pricing based on the distribution of $K^{D}(\xi)$ will provide significantly different results and this point should be carefully taken into consideration. Let us take a close look at $\lambda_{\alpha^{*}}$ (Figure 2) and $\lambda^{A,c}_{\alpha^{*}}$ (Figure 3) of the two exceptional ‘DAL’ and ‘KSS’. The density of both $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ in the case of ‘DAL’ is concentrated around 6 years, around the same levels of the other (six) companies while ‘KSS’ has concentration at much lower level. Taking into account that the $K^{D}(\xi)$ of these companies have similar concentration, $K^{D}(\xi)$ and $\lambda_{\alpha^{*}}$ (or $\lambda_{\alpha^{*}}^{A,c}$) provide different aspects of credit risk management.